# HippoRAG: Neurobiologically Inspired Long-Term Memory for Large Language Models

### v1

@@ Abstract
In order to thrive in hostile and ever-changing natural environments, mammalian brains evolved to store large amounts of knowledge about the world and continually integrate new information while avoiding catastrophic forgetting. Despite the impressive accomplishments, large language models (LLMs), even with retrieval-augmented generation (RAG), still struggle to efficiently and effectively integrate a large amount of new experiences after pre-training. In this work, we introduce HippoRAG, a novel retrieval framework inspired by the hippocampal indexing theory of human long-term memory to enable deeper and more efficient knowledge integration over new experiences. HippoRAG synergistically orchestrates LLMs, knowledge graphs, and the Personalized PageRank algorithm to mimic the different roles of neocortex and hippocampus in human memory. We compare HippoRAG with existing RAG methods on multi-hop question answering and show that our method outperforms the state-of-the-art methods remarkably, by up to $20$%. Single-step retrieval with HippoRAG achieves comparable or better performance than iterative retrieval like IRCoT while being $10$-$30$ times cheaper and $6$-$13$ times faster, and integrating HippoRAG into IRCoT brings further substantial gains. Finally, we show that our method can tackle new types of scenarios that are out of reach of existing methods. [^1] Code and data are available at https://github.com/OSU-NLP-Group/HippoRAG.
@@ 1 Introduction
Despite the progress of large language models (LLMs) in recent years, such a continuously updating long-term memory is still conspicuously absent from current AI systems. Due in part to its ease of use and the limitations of other techniques such as model editing [35], retrieval-augmented generation (RAG) has become the de facto solution for long-term memory in LLMs, allowing users to present new knowledge to a static model [28, 33, 50].
@@ 2.3 Detailed Methodology
Online Retrieval. During the retrieval process, we prompt $L$ using a 1-shot prompt to extract a set of named entities from a query $q$, our previously defined query named entities $C_{q}=\{c_{1},...,c_{n}\}$ (Stanford and Alzheimer’s in our Figure 4 example). These named entities $C_{q}$ from the query are then encoded by the same retrieval encoder $M$. Then, the previously defined query nodes are chosen as the set of nodes in $N$ with the highest cosine similarity to the query named entities $C_{q}$. More formally, query nodes are defined as $R_{q}=\{r_{1},...,r_{n}\}$ such that $r_{i}=e_{k}$ where $k=\operatorname*{arg\,max}_{j}cosine\_similarity(M(c_{i}),M(e_{j}))$, represented as the Stanford logo and the Alzheimer’s purple ribbon symbol in Figure 4
@@ 4 Results
Table 3: Multi-step retrieval performance. Combining HippoRAG with standard multi-step retrieval methods like IRCoT results in significant improvements on all three datasets.
@@ 5.1 What Makes HippoRAG Work?
HippoRAG $\mathbf{40.9}$ $\mathbf{51.9}$ $70.7$ $\mathbf{89.1}$ $\mathbf{60.5}$ $\mathbf{77.7}$ $\mathbf{57.4}$ $\mathbf{72.9}$
@@
Llama-3-8B-Instruct [1] $39.4$ $50.9$ $\mathbf{71.1}$ $88.0$ $59.4$ $76.4$ $56.6$ $71.8$
Llama-3-70B-Instruct [1] $36.3$ $45.9$ $56.9$ $68.8$ $51.6$ $66.1$ $48.3$ $60.3$
@@
OpenIE Alternatives. To determine if using GPT-3.5 is essential for building our KG, we replace it with an end-to-end OpenIE model REBEL [26] and an instruction-tuned LLM Llama-3 [1]. As shown in Table 5 row 2, building our KG using REBEL results in large performance drops, underscoring the importance of LLM flexibility. Specifically, GPT-3.5 produces twice as many triples as REBEL, indicating its bias against producing triples with general concepts and leaving many useful associations behind.
In terms of open-source LLMs, Table 5 (rows 3-4) shows that the performance of Llama-3 8B is comparable to GPT-3.5, although its 70B counterpart performs worse. This surprising behavior is due to this model’s production of ill-formatted outputs that result in the loss of around 20% of the passages, compared to about 4% for the 8B model and less than 1% for GPT-3.5. The strong performance of Llama-3 8B is encouraging because that offers a cheaper alternative for indexing over large corpora. The statistics for these OpenIE alternatives can be found in Appendix C.
@@ 5.2 HippoRAG’s Advantage: Single-Step Multi-Hop Retrieval
A major advantage of HippoRAG over conventional RAG methods in multi-hop QA is its ability to perform multi-hop retrieval in a single step. We demonstrate this by measuring the percentage of queries where all the supporting passages are retrieved successfully, a feat that can only be accomplished through successful multi-hop reasoning. Table 8 in Appendix D shows that the gap between our method and ColBERTv2, using the top-$5$ passages, increases even more from $3$% to $6$% on MuSiQue and from $20$% to $38$% on 2WikiMultiHopQA, suggesting that large improvements come from obtaining all supporting documents rather than achieving partially retrieval on more questions.
@@ 6 Related Work
Parametric Long-Term Memory. It is well-accepted, even among skeptical researchers, that the parameters of modern LLMs encode a remarkable amount of world knowledge [2, 10, 17, 21, 24, 31, 47, 62], which can be leveraged by an LLM in flexible and robust ways [64, 65, 74]. Nevertheless, our ability to update this vast knowledge store, an essential part of any long-term memory system, is still surprisingly limited. Although many techniques to update LLMs exist, such as standard fine-tuning, model unlearning and model editing [12, 37, 38, 39, 40, 76], it is clear that no methodology has emerged as a robust solution for continual learning in LLMs [20, 35, 78].
@@
Long Context as Long-Term Memory. Context lengths for both open and closed source LLMs have increased dramatically in the past year [9, 13, 16, 46, 51]. This scaling trend seems to indicate that future LLMs could perform long-term memory storage within massive context windows. However, the viability of this future remains largely uncertain given the many engineering hurdles involved and the apparent limitations of long-context LLMs, even within current context lengths [32, 34, 77].
LLMs & KGs. Combining the strengths of language models and knowledge graphs has been an active research direction for many years, both for augmenting LLMs with a KG in different ways [36, 63, 66] or augmenting KGs by either distilling knowledge from an LLM’s parametric knowledge [5, 67] or using them to parse text directly [6, 22, 75]. In an exceptionally comprehensive survey, Pan et al. [43] present a roadmap for this research direction and highlight the importance of work which synergizes these two important technologies [29, 57, 72, 80]. Like these works, HippoRAG is a strong and principled example of the synergy we must strike between these two technologies, combining the power of LLMs for knowledge graph construction with the strengths of structured knowledge and graph search for improved augmentation of an LLM’s capacities.
@@ 7 Conclusions & Limitations
Nevertheless, several limitations can be addressed in future work to enable HippoRAG to achieve this goal better. First, we note that all components of HippoRAG are currently used off-the-shelf without any extra training. There is therefore much room to improve our method’s practical viability by performing specific component fine-tuning. This is evident in the error analysis discussed in Appendix F that shows most errors made by our system are due to NER and OpenIE, which could benefit from direct fine-tuning. Given that the rest of the errors are graph search errors, also in Appendix F, we note that several avenues for improvements over simple PPR exist, such as allowing relations to guide graph traversal directly. Finally, and perhaps most importantly, HippoRAG’s scalability still calls for further validation. Although we show that Llama-3 could obtain similar performance to closed-source models and thus reduce costs considerably, we are yet to empirically prove the efficiency and efficacy of our synthetic hippocampal index as its size grows way beyond current benchmarks.
@@ Acknowledgments
  • Appendix D: Single-Step Multi-Hop: All-Recall Metric
@@
  • Appendix F: Error Analysis on MuSiQue
@@ Appendix B Dataset Comparison
Figure 6: Density of similarity scores of candidate passages (distractors and supporting passages) obtained by Contriever. The similarity score of HotpotQA distractors is not significantly larger than that of the least similar supporting passages, meaning that these distractors are not very effective.
@@ Appendix C Ablation Statistics
We use GPT-3.5 Turbo, REBEL [26] and Llama-3 (8B and 70B) [1] for OpenIE ablation experiments. As shown in Table 7, compared to both GPT-3.5 Turbo and both Llama models, REBEL generates around half the number of nodes and edges. This illustrates REBEL’s lack of flexibility in open information extraction when compared to using both open and closed-source LLMs. Meanwhile, both Llama-3 versions produce a similar amount of OpenIE triples than GPT-3.5 Turbo.
@@
GPT-3.5 Turbo [42] (Default) # of Unique Nodes ($N$) $91,729$ $42,694$ $82,157$
@@
REBEL [26] # of Unique Nodes ($N$) $36,653$ $22,146$ $30,426$
@@
Llama-3-8B-Instruct [1] # of Unique Nodes ($N$) $89,759$ $44,251$ $78,959$
# of Unique Edges ($E$) $25,868$ $8,497$ $20,325$
# of Unique Triples $117,127$ $54,347$ $101,822$
# of ColBERTv2 Synonym Edges ($E^{\prime}$) $186,568$ $92,913$ $157,446$
Llama-3-70B-Instruct [1] # of Unique Nodes ($N$) $83,647$ $38,800$ $73,958$
# of Unique Edges ($E$) $20,694$ $6,917$ $15,663$
# of Unique Triples $111,996$ $51,379$ $100,372$
# of ColBERTv2 Synonym Edges ($E^{\prime}$) $177,979$ $75,463$ $150,291$
Appendix D Single-Step Multi-Hop: All-Recall Metric
HippoRAG performs better on multi-hop QA benchmarks than standard RAG mainly because of its ability to perform multi-hop retrieval in a single step. Table 8 demonstrates this by measuring the percentage of queries where all supporting passages are retrieved successfully, something that can only be achieved through successful multi-hop reasoning. The fact that the gaps shown in this table are much larger than the gaps in Table 2 indicates that much of HippoRAG’s improvements are coming from the ability to retrieve all supporting documents not by performing better partial retrieval.
MuSiQue 2Wiki HotpotQA Average
AR@2 AR@5 AR@2 AR@5 AR@2 AR@5 AR@2 AR@5
ColBERTv2 [53] $6.8$ $16.1$ $25.1$ $37.1$ $33.3$ $59.0$ $21.7$ $37.4$
HippoRAG $10.2$ $22.4$ $45.4$ $75.7$ $33.8$ $57.9$ $29.8$ $52.0$
@@ F.1 Overview
We mainly analyze the errors by HippoRAG on the MuSiQue dataset. As shown in Table 10, these errors can be attributed to three types. The NER limitation is the main error type (nearly half), as NER may not extract enough information from the query for retrieval, e.g., for the question “When was one internet browser’s version of Windows 8 made accessible?”, only the phrase “Windows 8” is extracted. Nothing about “browser” or “accessibility” is extracted for subsequent graph search.
For OpenIE errors, we discuss some examples in §F.3. Despite the correct functioning of NER and OpenIE, the PPR algorithm sometimes fails to identify relevant passages in the graph correctly. For instance, consider the question “How many refugees emigrated to the European country where Huguenots felt a kinship for emigration?”. Despite the supporting passages being titled “Huguenots” as shown in the question, the PPR algorithm struggles due to multiple passages with similar topics within the corpus. Even when the term “Huguenots” is accurately extracted from both the question and the supporting passages, and the PPR algorithm initiates searches with the nodes labeled “European” and “Huguenots”, it incorrectly prioritizes other passages containing “European” and “Huguenots” that do not actually support the question. This misdirection occurs because the algorithm does not effectively discriminate between passages with similar topics but different relevance to the question’s context.
@@
Unfortunately, this design is also one of our method’s greatest limitations since ignoring contextual cues accounts for around $48$% of errors in a small-scale error analysis; more details can be found in Appendix F. This problem is more apparent in the second example since the concepts are general, making the context more important. Since the only concept tagged by HippoRAG is “protons”, it extracts passages related to “Uranium” and “nuclear weapons” while ColBERTv2 uses the context to extract more relevant passages associated with the discovery of atomic numbers.
@@
Table 12: Single-step retrieval performance. HippoRAG performs significantly better on MuSiQue and 2WikiMultiHopQA than all baselines and achieves comparable performance on the less challenging HotpotQA dataset.
@@
HippoRAG w/ Uncertainty-based Ensemble Contriever $\underline{42.3}$ $\underline{54.5}$ $71.3$ $87.2$ $60.6$ $79.1$ $\underline{58.1}$ $\underline{73.6}$
@@
To get a better trade-off between concepts and context, we introduce an ensembling setting where HippoRAG scores are ensembled with dense retrievers when our parahippocampal region shows uncertainty regarding the link between query and KG entities. This process represents instances when no hippocampal index was fully activated by the upstream parahippocampal signal and thus the neocortex must be relied on more strongly. We only use uncertainty-based ensembling if one of the query-KG entity scores $cosine\_similarity(M(c_{i}),M(e_{j}))$ is lower than a threshold $\theta$, for example, if there was no Stanford node in the KG and the closest node in the KG is something that has a cosine similarity lower than $\theta$ such as Stanford Medical Center. The final passage score for uncertainty-based ensembling is the average of the HippoRAG scores and standard passage retrieval using model $M$, both of which are first normalized into the 00 to $1$ over all passages.
When HippoRAG is ensembled with $M$ under “Uncertainty-based Ensembling”, it further improves on MuSiQue and outperforms our baselines in R@5 for HotpotQA, as shown in Table 12. When used in combination with IRCoT, as shown in Table 13, the ColBERTv2 ensemble outperforms all previous baselines in both R@2 and R@5 on HotpotQA. Although the simplicity of this approach is promising, more work needs to be done to solve this context-context tradeoff since simple ensembling does lower performance in some cases, especially for the 2WikiMultiHopQA dataset.
Table 13: Multi-step retrieval performance. Combining HippoRAG with standard multi-step retrieval methods like IRCoT results in significant improvements on all three datasets.
@@
IRCoT + HippoRAG w/ Uncertainty-based Ensemble Contriever $\underline{44.4}$ $\mathbf{58.5}$ $75.3$ $91.5$ $66.9$ $\underline{85.0}$ $\underline{62.2}$ $\mathbf{78.3}$

### v3

@@ Abstract
In order to thrive in hostile and ever-changing natural environments, mammalian brains evolved to store large amounts of knowledge about the world and continually integrate new information while avoiding catastrophic forgetting. Despite their impressive accomplishments, large language models (LLMs), even with retrieval-augmented generation (RAG), still struggle to efficiently and effectively integrate a large amount of new experiences after pre-training. In this work, we introduce HippoRAG, a novel retrieval framework inspired by the hippocampal indexing theory of human long-term memory to enable deeper and more efficient knowledge integration over new experiences. HippoRAG synergistically orchestrates LLMs, knowledge graphs, and the Personalized PageRank algorithm to mimic the different roles of neocortex and hippocampus in human memory. We compare HippoRAG with existing RAG methods on multi-hop question answering (QA) and show that our method outperforms the state-of-the-art methods remarkably, by up to $20$%. Single-step retrieval with HippoRAG achieves comparable or better performance than iterative retrieval like IRCoT while being $10$-$20$ times cheaper and $6$-$13$ times faster, and integrating HippoRAG into IRCoT brings further substantial gains. Finally, we show that our method can tackle new types of scenarios that are out of reach of existing methods. [^1] Code and data are available at https://github.com/OSU-NLP-Group/HippoRAG.
@@ 1 Introduction
Despite the progress of large language models (LLMs) in recent years, such a continuously updating long-term memory is still conspicuously absent from current AI systems. Due in part to its ease of use and the limitations of other techniques such as model editing [46], retrieval-augmented generation (RAG) has become the de facto solution for long-term memory in LLMs, allowing users to present new knowledge to a static model [36, 42, 66, 87].
@@ 2.3 Detailed Methodology
Online Retrieval. During the retrieval process, we prompt $L$ using a 1-shot prompt to extract a set of named entities from a query $q$, our previously defined query named entities $C_{q}=\{c_{1},...,c_{n}\}$ (Stanford and Alzheimer’s in our Figure 4 example). These named entities $C_{q}$ from the query are then encoded by the same retrieval encoder $M$. Then, the previously defined query nodes are chosen as the set of nodes in $N$ with the highest cosine similarity to the query named entities $C_{q}$. More formally, query nodes are defined as $R_{q}=\{r_{1},...,r_{n}\}$ such that $r_{i}=e_{k}$ where $k=\operatorname*{arg\,max}_{j}cosine\_similarity(M(c_{i}),M(e_{j}))$, represented as the Stanford logo and the Alzheimer’s purple ribbon symbol in Figure 4.
@@ 4 Results
Table 3: Multi-step retrieval performance. Combining HippoRAG with standard multi-step retrieval methods like IRCoT results in strong complementary improvements on all three datasets.
@@ 5.1 What Makes HippoRAG Work?
HippoRAG $40.9$ $51.9$ $\mathbf{70.7}$ $\mathbf{89.1}$ $60.5$ $77.7$ $\mathbf{57.4}$ $\mathbf{72.9}$
@@
Llama-3.1-8B-Instruct [1] $40.8$ $51.9$ $62.5$ $77.5$ $59.9$ $75.1$ $54.4$ $67.8$
Llama-3.1-70B-Instruct [1] $\mathbf{41.8}$ $\mathbf{53.7}$ $68.8$ $85.3$ $\mathbf{60.8}$ $\mathbf{78.6}$ $57.1$ $72.5$
@@
OpenIE Alternatives. To determine if using a closed model like GPT-3.5 is essential to retain our performance improvements, we replace it with an end-to-end OpenIE model REBEL [34] as well as the 8B and 70B instruction-tuned versions of Llama-3.1, a class of strong open-weight LLMs [1]. As shown in Table 5 row 2, building our KG using REBEL results in large performance drops, underscoring the importance of LLM flexibility. Specifically, GPT-3.5 produces twice as many triples as REBEL, indicating its bias against producing triples with general concepts and leaving many useful associations behind.
In terms of open-weight LLMs, Table 5 (rows 3-4) shows that the performance of Llama-3.1-8B is competitive with GPT-3.5 in all datasets except for 2Wiki, where performance drops substantially. Nevertheless, the stronger 70B counterpart outperforms GPT-3.5 in two out of three datasets and is still competitive in 2Wiki. The strong performance of Llama-3.1-70B and the comparable performance of even the 8B model is encouraging since it offers a cheaper alternative for indexing over large corpora. The graph statistics for these OpenIE alternatives can be found in Appendix C.
To understand the relationship between OpenIE and retrieval performance more deeply, we extract 239 gold triples from 20 examples from the MuSiQue training set. We then perform a small-scale intrinsic evaluation using the CaRB [6] framework for OpenIE. We find that both Llama-3.1-Instruct models underperform GPT-3.5 slightly on this intrinsic evaluation but all LLMs vastly outperform REBEL. More details about this evaluation experiments can be found in Appendix D.
@@ 5.2 HippoRAG’s Advantage: Single-Step Multi-Hop Retrieval
A major advantage of HippoRAG over conventional RAG methods in multi-hop QA is its ability to perform multi-hop retrieval in a single step. We demonstrate this by measuring the percentage of queries where all the supporting passages are retrieved successfully, a feat that can only be accomplished through successful multi-hop reasoning. Table 6 below shows that the gap between our method and ColBERTv2, using the top-$5$ passages, increases even more from $3$% to $6$% on MuSiQue and from $20$% to $38$% on 2WikiMultiHopQA, suggesting that large improvements come from obtaining all supporting documents rather than achieving partially retrieval on more questions.
MuSiQue 2Wiki HotpotQA Average
AR@2 AR@5 AR@2 AR@5 AR@2 AR@5 AR@2 AR@5
ColBERTv2 [70] $6.8$ $16.1$ $25.1$ $37.1$ $33.3$ $59.0$ $21.7$ $37.4$
HippoRAG $10.2$ $22.4$ $45.4$ $75.7$ $33.8$ $57.9$ $29.8$ $52.0$
@@ 6 Related Work
6.1 LLM Long-Term Memory
Parametric Long-Term Memory. It is well-accepted, even among skeptical researchers, that the parameters of modern LLMs encode a remarkable amount of world knowledge [2, 12, 23, 28, 31, 39, 62, 79], which can be leveraged by an LLM in flexible and robust ways [81, 83, 93]. Nevertheless, our ability to update this vast knowledge store, an essential part of any long-term memory system, is still surprisingly limited. Although many techniques to update LLMs exist, such as standard fine-tuning, model editing [15, 49, 50, 51, 52, 95] and even external parametric memory modules inspired by human memory [58, 82, 32], no methodology has yet to emerge as a robust solution for continual learning in LLMs [26, 46, 97].
@@ 6.1 LLM Long-Term Memory
Long Context as Long-Term Memory. Context lengths for both open and closed source LLMs have increased dramatically in the past year [11, 17, 22, 61, 68]. This scaling trend seems to indicate that future LLMs could perform long-term memory storage within massive context windows. However, the viability of this future remains largely uncertain given the many engineering hurdles involved and the apparent limitations of long-context LLMs, even within current context lengths [41, 45, 96, 21].
6.2 Multi-Hop QA & Graphs
Many previous works have also tackled multi-hop QA using graph structures. These efforts can be broadly divided in two major categories: 1) graph-augmented reading comprehension, where a graph is extracted from retrieved documents and used to improve a model’s reasoning process and 2) graph-augmented retrieval, where models find relevant documents by traversing a graph structure.
Graph-Augmented Reading Comprehension. Earlier works in this category are mainly supervised methods which mix signal from a hyperlink or co-occurrence graph with a language model through a graph neural network (GNN) [20, 67, 65]. More recent works use LLMs and introduce knowledge graph triples directly into the LLM prompt [57, 43, 47]. Although these works share HippoRAG’s use of graphs for multi-hop QA, their generation-based improvements are fully complementary to HippoRAG’s, which are solely based on improved retrieval.
Graph-Augmented Retrieval. In this second category, previous work trains a re-ranking module which can traverse a graph made using Wikipedia hyperlinks [16, 100, 54, 14, 4, 44]. HippoRAG, in contrast, builds a KG from scratch using LLMs and performs multi-hop retrieval without any supervision, making it much more adaptable.
6.3 LLMs & KGs
Combining the strengths of language models and knowledge graphs has been an active research direction for many years, both for augmenting LLMs with a KG in different ways [48, 80, 84] or augmenting KGs by either distilling knowledge from an LLM’s parametric knowledge [7, 85] or using them to parse text directly [8, 29, 94]. In an exceptionally comprehensive survey, Pan et al. [56] present a roadmap for this research direction and highlight the importance of work which synergizes these two important technologies [37, 74, 27, 91, 99]. Like these works, HippoRAG shows the potential for synergy between these two technologies, combining the knowledge graph construction abilities of LLMs with the retrieval advantages of structured knowledge for more effective RAG.
@@ 7 Conclusions & Limitations
Nevertheless, several limitations can be addressed in future work to enable HippoRAG to achieve this goal better. First, we note that all components of HippoRAG are currently used off-the-shelf without any extra training. There is therefore much room to improve our method’s practical viability by performing specific component fine-tuning. This is evident in the error analysis discussed in Appendix F, which shows most errors made by our system are due to NER and OpenIE and thus could benefit from direct fine-tuning. Given that the rest of the errors are graph search errors, also in Appendix F, we note that several avenues for improvements over simple PPR exist, such as allowing relations to guide graph traversal directly. Additionally, as shown in Appendix F.4, more work must be done to improve the consistency of OpenIE in longer compared to shorter documents. Finally, and perhaps most importantly, HippoRAG’s scalability still calls for further validation. Although we show that Llama-3.1 could obtain similar performance to closed-source models and thus reduce costs considerably, we are yet to empirically prove the efficiency and efficacy of our synthetic hippocampal index as its size grows way beyond current benchmarks.
@@ Acknowledgments
  • Appendix D: Intrinsic OpenIE Evaluation
@@
  • Appendix F: Error Analysis
@@ Appendix B Dataset Comparison
Figure 6: Density of similarity scores of candidate passages (distractors and supporting passages) obtained by Contriever. The similarity score of HotpotQA distractors is not substantially larger than that of the least similar supporting passages, meaning that these distractors are not very effective.
@@ Appendix C Ablation Statistics
We use GPT-3.5 Turbo, REBEL [34] and Llama-3.1 (8B and 70B) [1] for OpenIE ablation experiments. As shown in Table 8, compared to both GPT-3.5 Turbo and both Llama models, REBEL generates around half the number of nodes and edges. This illustrates REBEL’s lack of flexibility in open information extraction when compared to using both open and closed-source LLMs. Meanwhile, both Llama-3.1 versions produce a similar amount of OpenIE triples than GPT-3.5 Turbo.
@@
GPT-3.5 Turbo (1106) [55] (Default) # of Unique Nodes ($N$) $91,729$ $42,694$ $82,157$
@@
REBEL-large [34] # of Unique Nodes ($N$) $36,653$ $22,146$ $30,426$
@@
Llama-3.1-8B-Instruct [1] # of Unique Nodes ($N$) $86,864$ $37,875$ $76,311$
# of Unique Edges ($E$) $22,807$ $6,729$ $18,109$
# of Unique Triples $118,430$ $47,420$ $104,981$
# of ColBERTv2 Synonym Edges ($E^{\prime}$) $155,889$ $72,963$ $139,181$
Llama-3.1-70B-Instruct [1] # of Unique Nodes ($N$) $80,634$ $39,845$ $70,304$
# of Unique Edges ($E$) $22,120$ $6,996$ $16,404$
# of Unique Triples $120,514$ $55,940$ $105,281$
# of ColBERTv2 Synonym Edges ($E^{\prime}$) $140,328$ $69,125$ $119,948$
Appendix D Intrinsic OpenIE Evaluation
In order to better understand how OpenIE and retrieval interact, we extracted gold triples from 20 documents from the MuSiQue training dataset. In total, we extracted 239 gold triples. From the results in Table 9, we first note that there is a massive difference between end-to-end information extraction systems like REBEL and LLMs. Additionally, we note that there is some correlation better OpenIE and retrieval performance, given that the 8B Llama-3.1-Instruct version performs worse that its 70B counterpart in both retrieval and intrinsic metrics. More specifically, we see that this larger model only provides intrinsic improvements in the recall metric, which seems specially important in improving retrieval performance. Finally, we note that this evaluation is not perfectly correlated with retrieval performance, since GPT-3.5’s intrinsic performance is much stronger than Llama-3.1-70B-Instruct while its retrieval score is only slightly higher.
Table 9: Intrinsic OpenIE evaluation using the CaRB [6] framework on 20 annotated passages.
AUC Precision Recall F1
GPT-3.5 Turbo (1106) [55] (Default) $46.5$ $68.4$ $55.2$ $61.1$
Llama-3.1-8B-Instruct [1] $40.0$ $66.4$ $48.1$ $55.8$
Llama-3.1-70B-Instruct [1] $42.3$ $66.3$ $50.9$ $57.6$
REBEL [34] $1.0$ $8.0$ $1.8$ $2.9$
@@ F.1 Overview
In this section, we provide a detailed error analysis of 100 errors made by HippoRAG on the MuSiQue dataset. As shown in Table 11, these errors can be categorized into three main types: NER, OpenIE and PPR.
The main error type, with nearly half of all error examples, is due to limitations of our NER based design. As further discussed in §F.2, our NER design does not extract enough information from the query for retrieval. For example, in the question “When was one internet browser’s version of Windows 8 made accessible?”, only the phrase “Windows 8” is extracted, leaving any signal about “browsers” or “accessibility” behind for the subsequent graph search. OpenIE errors, the second most common, are discussed in more detail in §F.3.
We define the third error category as cases where both NER and OpenIE are functioning properly but the PPR algorithm is still unable to identify relevant subgraphs, often due to confounding signals. For instance, consider the query “How many refugees emigrated to the European country where Huguenots felt a kinship for emigration?”. Despite the term “Huguenots” being accurately extracted from both the question and the supporting passages, and the PPR algorithm initiating with the nodes labeled “European” and “Huguenots”, the PPR algorithm struggles to find the appropriate subgraphs around them that define the most related passage. This occurs when multiple passages exist in the corpus that discuss very similar topics since the PPR algorithm is not able to leverage query context directly.
@@
Unfortunately, this design is also one of our method’s greatest limitations since ignoring contextual cues accounts for around $48$% of errors in our small-scale error analysis. This problem is more apparent in the second example since the concepts are general, making the context more important. Since the only concept tagged by HippoRAG is “protons”, it extracts passages related to “Uranium” and “nuclear weapons” while ColBERTv2 uses the context to extract more relevant passages associated with the discovery of atomic numbers.
@@
Table 13: Single-step retrieval performance. HippoRAG performs substantially better on MuSiQue and 2WikiMultiHopQA than all baselines and achieves comparable performance on the less challenging HotpotQA dataset.
@@
HippoRAG w/ Uncertainty Ensemble Contriever $\underline{42.3}$ $\underline{54.5}$ $71.3$ $87.2$ $60.6$ $79.1$ $\underline{58.1}$ $\underline{73.6}$
@@
To get a better trade-off between concepts and context, we introduce an ensembling setting where HippoRAG scores are ensembled with dense retrievers when our parahippocampal region shows uncertainty regarding the link between query and KG entities. This process represents instances when no hippocampal index was fully activated by the upstream parahippocampal signal and thus the neocortex must be relied on more strongly. We only use uncertainty ensembling if one of the query-KG entity scores $cosine\_similarity(M(c_{i}),M(e_{j}))$ is lower than a threshold $\theta$, for example, if there was no Stanford node in the KG and the closest node in the KG is something that has a cosine similarity lower than $\theta$ such as Stanford Medical Center. The final passage score for uncertainty ensembling is the average of the HippoRAG scores and standard passage retrieval using model $M$, both of which are first normalized into the 00 to $1$ over all passages.
When HippoRAG is ensembled with $M$ under “Uncertainty Ensemble”, it further improves on MuSiQue and outperforms our baselines in R@5 for HotpotQA, as shown in Table 13. When used in combination with IRCoT, as shown in Table 14, the ColBERTv2 ensemble outperforms all previous baselines in both R@2 and R@5 on HotpotQA. Although the simplicity of this approach is promising, more work needs to be done to solve this context-context tradeoff since simple ensembling does lower performance in some cases, especially for the 2WikiMultiHopQA dataset.
Table 14: Multi-step retrieval performance. Combining HippoRAG with standard multi-step retrieval methods like IRCoT results in substantial improvements on all three datasets.
@@
IRCoT + HippoRAG w/ Uncertainty Ensemble Contriever $\underline{44.4}$ $\mathbf{58.5}$ $75.3$ $91.5$ $66.9$ $\underline{85.0}$ $\underline{62.2}$ $\mathbf{78.3}$